OpenStudy (anonymous):

consider the piecewise h(x) = -2/3(x+1) x<-1 -4x-4 -1≤ x≤ 0 x-8 1

OpenStudy (anonymous):

http://awesomescreenshot.com/0303lr6g9d

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@ganeshie8

jimthompson5910 (jim_thompson5910):

So they want you to find \[\Large h^{-1}(3)\] right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

if you can show me how to do one i can do the rest

jimthompson5910 (jim_thompson5910):

that's effectively asking: what is the value of x when h(x) = 3

jimthompson5910 (jim_thompson5910):

put another way, "what is the x value for the point (x,3)" this point lies on the piecewise function

jimthompson5910 (jim_thompson5910):

so you'd go to the graph and find where the point with a y coordinate of 3 is it is on the left most piece, the red piece so you'd solve -2/3(x+1) = 3 for x

OpenStudy (anonymous):

so are the answers 7/2, -1, and 11?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

also, is h(x) invertible?

jimthompson5910 (jim_thompson5910):

how are you getting 7/2 ?

OpenStudy (anonymous):

i solved for x. sorry i meant -3, 0, and 3 respectively so i got 7/2, -1, and 11

jimthompson5910 (jim_thompson5910):

-2/3(x+1) = 3 -2(x+1) = 3*3 -2(x+1) = 9 -2x-2 = 9 -2x = 9+2 -2x = 11 x = -11/2

jimthompson5910 (jim_thompson5910):

So that effectively says: h(-11/2) = 3

jimthompson5910 (jim_thompson5910):

you are correct in stating that h(-1) = 0

jimthompson5910 (jim_thompson5910):

h(11) is not defined because the function only goes up to x = 4, but it doesn't include 4 itself. This is the furthest you can go to the right.

OpenStudy (anonymous):

the question says evaluate h^-1(-3), h^-1(0), h^-1(3) so it has to be h^-1(3)=7/2, not h^-1(7/2)=3

OpenStudy (anonymous):

h^-1(-3)=7/2, not h^-1(7/2)=-3

jimthompson5910 (jim_thompson5910):

oh we're talking about two different things I'm saying that because h(-11/2) = 3, this means h^-1(3) = -11/2

OpenStudy (anonymous):

but h^-1(3) = 7/2, not -11/2

jimthompson5910 (jim_thompson5910):

7/2 = 3.5 isn't in the domain of h(x) though

OpenStudy (anonymous):

I'm really confused so when i evaluated h, i got h(-3)=4/3, h(0)=-4, and h(3)=-5

OpenStudy (anonymous):

so i have to find the inverse of h(-3), h(0), and h(3)

jimthompson5910 (jim_thompson5910):

when you find the inverse of a function, like y = x+2, you are simply solving for y to get x = y - 2

OpenStudy (anonymous):

but when i find the inverse of h(-3), i should get 4/3, but i get 7/2

jimthompson5910 (jim_thompson5910):

when they ask: "find the inverse of f^-1(2)" they mean plug in y = 2 and solve for x x = y-2 x = 2-2 x = 0 so that tells us f(0) = 2

jimthompson5910 (jim_thompson5910):

so the book is saying the answer is 4/3?

OpenStudy (anonymous):

there is no answer key

OpenStudy (anonymous):

but my graphic and algebraic answer is different

jimthompson5910 (jim_thompson5910):

ok let's do h^-1(3)

jimthompson5910 (jim_thompson5910):

step 1) locate 3 on the y axis step 2) draw a horizontal line through 3 on the y axis step 3) find out where this horizontal line intersects with the graph of h(x). This happens at x = -11/2 and you can determine that by solving -2/3(x+1) = 3 Those steps tell us that h(-11/2) = 3

jimthompson5910 (jim_thompson5910):

also, -11/2 = -6.5 so you should be exactly between -5 and -6 on the x axis (where the horizontal line and h(x) intersect)

OpenStudy (anonymous):

but h^-1(3) is supposed to be h(x)=x-8

OpenStudy (anonymous):

so its h^-1(x)=3+8=11

jimthompson5910 (jim_thompson5910):

but that green piece is nowhere near y = 3

OpenStudy (anonymous):

but according to h(x)=x-8, 1<x<4 and 1<3<4

jimthompson5910 (jim_thompson5910):

x isn't 3, y is

jimthompson5910 (jim_thompson5910):

if you evaluated h(3), then yes, you'd say x = 3 is between 1 and 4 (1 < x < 4 ---> 1 < 3 < 4) so h(3) = 3-8 = -5

jimthompson5910 (jim_thompson5910):

but that's not what they're asking

OpenStudy (anonymous):

wait so hold on: for h(x)=-2/3(x+1), x=-2/3y-2/3, y=-3/2x-1

jimthompson5910 (jim_thompson5910):

yes, then you would plug in x = 3

OpenStudy (anonymous):

okay so h^-1(3)=11/2

OpenStudy (anonymous):

-11/2

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

okay then i get h(x)=-4x-4 so h^-1(x)=-1/4x-1

OpenStudy (anonymous):

so do i plug in 0

OpenStudy (anonymous):

and then h^-1(0)=-1

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

so, h^-1(x)=x+8, so h^-1(-3)=5

jimthompson5910 (jim_thompson5910):

notice how 5 isn't in the domain though

OpenStudy (anonymous):

yea

jimthompson5910 (jim_thompson5910):

so h^-1(-3) is undefined

jimthompson5910 (jim_thompson5910):

it would be h^-1(-3) = 5 if you extended the domain out to x = 5

OpenStudy (anonymous):

so is h(x) invertible

jimthompson5910 (jim_thompson5910):

yes because it is one-to-one (it passes the horizontal line test)